OpenStudy (simoner):

Estimate the sum from n=1 to infinity of (2n+1)^(−9) correct to five decimal places.

OpenStudy (simoner):

\[\sum_{n=1}^{\infty}(2n+1)^{-9}\]

OpenStudy (perl):

The remainder (R) of a series that converges by the integral test of the first n terms can be estimated as follows: $$ \Large \int_{n+1}^{\infty}f(x) dx \leq R \leq \int_{n}^{\infty} f(x) dx $$

OpenStudy (perl):

Since we require 5 digit accuracy, we should aim for R < 0.00001. I'll use 0.000005.

OpenStudy (perl):

$$ \Large{ \int_{n+1}^{\infty}f(x) dx \leq R_n \leq \int_{n}^{\infty} \frac{1}{(2x+1)^9} dx \leq 0.000005 \\~\\ \therefore \\ ~\\ \left[ \frac{-1}{(16(2x+1)^8)} \right]_{n}^{\infty}\leq 0.000005 \\ ~\\ \therefore \frac{-1}{(16(2 \cdot \infty +1)^8)} - \frac{-1}{(16(2n+1)^8)} \leq 0.000005 \\ ~\\ \therefore \\ 0 + \frac{1}{(16(2n+1)^8)} \leq 0.000005 \\ ~\\ \therefore \\ \frac{1}{(16(2n+1)^8)} \leq 0.000005 \\ ~\\ \therefore \\ \frac{1}{(16(2n+1)^8)} \leq \frac{5}{1000000} } $$

OpenStudy (simoner):

ok i got\[\frac{ 1 }{ (2n+1)^{8} }\le \frac{ 1 }{ 12500 }\] \[(2n+1)^{8}\ge12500\] \[2n+1\ge3.251725\] \[2n \ge2.251725\] \[n \ge1.125862\]

OpenStudy (simoner):

can someone please help me with what to do next

OpenStudy (phi):

If all of that is correct, then you want to sum up the terms from n=1 to n=2

OpenStudy (simoner):

okay so i use \[\frac{ 1 }{ (2(1)+1)^{8} }+\frac{ 1 }{ (2(2)+1)^{8} }\]

OpenStudy (phi):

I think the original series has 9 not 8 as the exponent

OpenStudy (simoner):

oh yeah so that would equal 0.0000513

OpenStudy (simoner):

?

OpenStudy (simoner):

do i do anything after that?

OpenStudy (phi):

that looks about right if you add the next term you will notice the 5 does not change

OpenStudy (simoner):

but when i typed this in the answer was wrong

OpenStudy (simoner):

someone please help?

OpenStudy (amistre64):

what is the value when n=1?

OpenStudy (simoner):

1/19683

OpenStudy (amistre64):

yeah thats not in decimal form is it ....

OpenStudy (simoner):

0.0000508

OpenStudy (amistre64):

now compare this to what you got when n=2, do you see that they are the same for the first 5 decies?

OpenStudy (simoner):

yes

OpenStudy (amistre64):

then n=2 is not the answer we are looking for

OpenStudy (simoner):

but when i typed this in it said the answer was wrong

OpenStudy (amistre64):

0.00005 is correct to 5 decimal places

OpenStudy (amistre64):

im not sure how your program is wanting the answer formated. but im pretty sure that the estimation is: .00005

OpenStudy (amistre64):

can you screenshot it?

OpenStudy (amistre64):

yeah, you want it accurate to 5 decimal places, you arent putting in exactly 5 decimal places are you?

OpenStudy (amistre64):

0.12345 0.00005

OpenStudy (anonymous):

i do you need help still